1. A force of 40.0 N is applied to a square disk of side length 20.0 cm. What is the pressure exerted on the disk?

2. What volume is occupied by 1.00 mole of an ideal gas at 2.00 atm and a temperature of 20.0°C?

thank you in advance!
Answer #1

pressure on the square disk is force exerted per unit area = 40/( 0.2 x 0.2) = 1000 Pa

2)

ideal gas follows PV =nRT so V = 12.02135 L

here P is pressure of gas = 2 atm

V is volume of gas = ?

n is number of moles=1

T is temperature of gas = 20+273 = 293 K

R is gas constant = 0.082057 L atm K-1mol-1

please rate it up thanks :)